OpenStudy (anonymous):

describe the transformation of the parent function: g(x)=f(x)-4

OpenStudy (anonymous):

If a number is added/subtracted outside of a function (in your case f(x)), it means that is has moved up or down that many steps in regard to the parent function. So y=x^2+5 has moved 5 steps up above y=x^2